Banana Republic, the currency drain ratio is 0.45 of deposits and the reserve ratio is 0.05.
The country's neighbour, the Kingdom of Coconut, has a quantity of money of C$ 100 billion
(newly created money), the currency drain ratio is 40% of deposits and the reserve ratio is 10%
-calculate money multipier in banana Republic use increase of $100 000 as money base, how much money will be created?
-calculate the money base in the Kingdom of Coconut
1
Expert's answer
2019-10-21T08:17:46-0400

The money multipier in banana Republic is:

mm = 1/rr = 1/0.05 = 20.

In this case 100,000×20 = $2,000,000 will be created.

Deposits = Quantity of money×currency drain ratio = 100 billion×0.4 = $40 billion.

Money base = Currency + Deposits = 100 + 40 = $140 billion.